please provide an explanation for both questions! Thank You!9. Consider the pedigree to the right, displaying the inheritance of a rare recessive autosomal disease which is fully penetr
Answer #1

9) we have to find the probability that IV-1 is affected, for IV-1 to be affected the parents of IV-1 ( III-1 and III-2 ) has to be carriers, for III-1 to a carrier her father has to be a carrier and for III-2 to be a carrier his mother has to be a carrier.

the probability that II-2 is a carrier=?

II-2 and II-4 have an affected sister, but their parents are normal this shows the parents of II-2 and II-4 are carriers.

let the alleles be D- dominant normal, d- recessive

Dd * Dd

D d
D DD Dd
d Dd dd

probability of getting phenotypically normal carrier progenies= number of Dd/total number of progeny with D allele=2/3

probability of getting dd progenies for Dd parents= number of dd/total number=1/4

II-1 and II-5 are married to the family and they don`t have the disease so they are homozygous DD,

DD * Dd

D d
D DD Dd

probability of getting Dd progeny= number of Dd/total number=1/2

the probability that III-1 is a carrier= probability that II-2 is a carrier \times probability of getting Dd progenies for DD and Dd

= probability of getting phenotypically normal carrier progenies for Dd parents \times probability of getting Dd progenies for DD and Dd

= 2/3\times1/2

= 1/3

probability that III-2 is a carrier = probability that II-4 is a carrier \times probability of getting Dd progenies for DD and Dd

= probability of getting phenotypically normal carrier progenies for Dd parents \times probability of getting Dd progenies for DD and Dd

= 2/3\times1/2

= 1/3

probability that IV-1 is affected= probability that III-2 is a carrier \times probability that III-1 is a carrier \times probability of getting dd progenies for Dd parents

= 1/3\times1/3\times1/4

=1/36

so the answer is d) 1/36

10) the affected child has 2 copies of X chromosome and one Y chromosome, G6PD gene is located on the X chromosome. the phenotype of the affected child is A, so the child has inherited the X chromosome from mother only, ( if the child has inherited X chromosome from father it should have produced B phenotype).

mother has 2 X chromosomes one with A allele and the other with B allele, the homologous chromosomes are separated during meiosis I, if non-disjunction occurred in meiosis I both X chromosomes would have moved to one cell so the child would have got both A and B allele.

sister chromatids are separated during meiosis II, here the sister chromatids of the X chromosome with the A allele failed to separate in meiosis II, so one gamete has 2 X chromosomes with A allele which was fertilized with sperm containing Y chromosome so the genotype of the child is XXY.

here non-disjunction occurred in mother in meiosis II.

so the answer is b) mother, meiosis 2.